I have tried for an hour now, and I do not understand these equations.. For instance; how do you sovle this?
x + 2y = 13
2x - 3y = -9

Answer by jim_thompson5910(35256)

Multiply EVERY term in equation 1 by -2 to get

x + 2y = 13

-2x -4y = -26


We now have this system

-2x - 4y = -26
2x - 3y = -9


Add the equations to get


-2x - 4y = -26
2x - 3y = -9
-------------
0x - 7y = -35

So -7y = -35 which means y = 5 (divide both sides by -7)

Once you have y = 5, you use this to find x. Pick any equation with x,y in it and plug in y = 5 to solve for x. I'll let you do this part.